Two point charges are on the y-axis. A 4.50uC charge is located at y=1.25 cm, and a -2.24uC charge is located at y=-1.80 cm. Find the total electric potential at (a) the origin and (b) the point whose coordinates are (1.50 cm, 0).
1
Expert's answer
2020-11-12T09:37:16-0500

(a) The total electric potential produced by a two point charges is the sum of the electric potential produced by a 4.50uC charge and the electric potential produced by a -2.24uC charge:


"V_{tot}=V_1+V_2,""V_{tot}=k(\dfrac{Q_1}{r_1}+\dfrac{Q_2}{r_2})=k(\dfrac{Q_1}{y_1}+\dfrac{Q_2}{y_2}),"

here, "k=8.99\cdot10^9\ N\dfrac{m^2}{C^2}" is the Coulomb's constant, "Q_1=4.5\cdot10^{-6}\ C" is the first point charge, "Q_2=-2.24\cdot10^{-6}\ C" is the second point charge, "r_1=y_1=1.25\cdot10^{-2}\ m" is the distance from the first point charge to the origin, "r_2=y_2=1.8\cdot10^{-2}\ m" is the distance from the second point charge to the origin.

Then, we can calculate the total electric potential at the origin:

"V_{tot}=8.99\cdot10^9\ N\dfrac{m^2}{C^2}\cdot(\dfrac{4.5\cdot10^{-6}\ C}{1.25\cdot10^{-2}\ m}+\dfrac{(-2.24\cdot10^{-6}\ C)}{1.8\cdot10^{-2}\ m})=2.12\cdot10^6\ V."

(b) Since the point located at "x-"axis, we can find the distances "r_1" and "r_2" from the Pythagorean theorem:


"r_1=\sqrt{x^2+y_1^2}=\sqrt{(1.5\cdot10^{-2})^2+(1.25\cdot10^{-2})^2}=0.019\ m,""r_2=\sqrt{x^2+y_2^2}=\sqrt{(1.5\cdot10^{-2})^2+(1.8\cdot10^{-2})^2}=0.023\ m."

Then, we can calculate the total electric potential at the point whose coordinates are (1.50 cm, 0):


"V_{tot}=k(\dfrac{Q_1}{r_1}+\dfrac{Q_2}{r_2}),"

"V_{tot}=8.99\cdot10^9\ N\dfrac{m^2}{C^2}\cdot(\dfrac{4.5\cdot10^{-6}\ C}{0.019\ m}+\dfrac{(-2.24\cdot10^{-6}\ C)}{0.023\ m})=1.25\cdot10^6\ V."

Answer:

(a) "V_{tot}=2.12\cdot10^6\ V."

(b) "V_{tot}=1.25\cdot10^6\ V."